40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Feb 2 09:52:19 2021, aaron, Computing, DAQ, cymac to Debian buster 
    Reply  Wed Feb 3 08:23:12 2021, Chris, Computing, DAQ, cymac to Debian buster 
       Reply  Tue Feb 9 16:33:31 2021, aaron, Computing, DAQ, cymac to Debian buster 
          Reply  Thu Feb 11 18:51:39 2021, aaron, Computing, DAQ, cymac to Debian buster Screen_Shot_2021-02-11_at_19.27.21.png
             Reply  Mon Feb 15 18:25:19 2021, Chris, Computing, DAQ, cymac to Debian buster 
                Reply  Tue Feb 16 14:29:04 2021, aaron, Computing, DAQ, cymac to Debian buster 55B5D301-4243-45DA-9B03-45EF87B2A684.jpeg4C7FE7DC-E8A0-4770-9A49-14EB491F76E2.jpeg13D67AB6-5A77-4C94-A2A4-FE7FDF065B01.jpegC84B3054-269B-4969-817F-1429BB671C7A.jpeg
                   Reply  Thu Feb 18 16:16:38 2021, aaron, Computing, DAQ, cymac to Debian buster 15A01777-A1C5-4710-8180-28C8B68A07AE.jpegF90CD5EF-1A26-4C4A-9C84-046E49642545.jpeg
                      Reply  Fri Feb 19 14:43:09 2021, aaron, Computing, DAQ, AI chassis diagnosis 
                   Reply  Tue Mar 9 14:03:12 2021, aaron, Computing, DAQ, oma model 
Message ID: 2661     Entry time: Fri Feb 19 14:43:09 2021     In reply to: 2654
Author: aaron 
Type: Computing 
Category: DAQ 
Subject: AI chassis diagnosis 

Opened up the AI chassis

  • Interface card with pin mapping is D070101-v1. AI board is D070081-v2.
  • Rana and I identified that op amp U3A needed replacement on channel 1 only; all other channels were functioning normally
  • I removed the front panel and flipped over the offending circuit board ("A). I powered on the board to identify the correct op amp, but something was shorting the negative supply. I disconnected the flipped board from power, and the supply was still shorted -- this was happening somewhere on the board for channels 9-16 ("B"), which I hadn't touched. I disconnected power from board B, wiggled the connectors, and reconnected it -- power was restored. 
  • I again flipped over board A, and this time was able to identify the offending op amp by testing its input and output pins, and nearby resistors. I replaced the op amp by 
    1. apply tack flux SMD291 to all pins of IC
    2. melt chip quik alloy on pins, and with the alloy molten lift off the IC
    3. Clean off the brittle chip quik with a flux coated solder wick. Wipe away the remaining flux with a kim wipe and isopropyl
    4. Solder on the new surface mount op amp. It is an AD8672ARZ, which we had in the electronics shop.
  • Realized I had replaced the wrong op amp (U2, not U3). I hadn't realized the AD8672 is a dual op amp, so misunderstood the pin mapping from before. Fortunately, the swap was good and the changed op amp is still working.
  • After the swap, the negative supply short returned. Again, it was on board B, not the board I had been working on. After wiggling cables again, taking apart the connector and confirming no stray wires and good connections, the short went away and power was restored. Pretty confused by this. I might have just had the current limit too low? Could not reproduce the problem, and didn't do anything that should have fixed a short. 
  • After replacing the correct op amp, channel 1 no longer has an offset. yes

I meant to take a photo while still opened up, so I left the DAC to AI unplugged to remind myself to get to that next time; should also post some transfer functions to confirm everything is functioning. 

Update: photo is now on the west bridge wiki and ligo.wbridge google drive

ELOG V3.1.3-